What is the freezing point of a solution containing 17.98 g diphenylether (C12H10O), a nonelectrolyte, and 291.7 g acetic acid (CH3COOH)? The normal freezing point of acetic acid is 16.60 °C and Kfp for acetic acid is 3.90 °C/m.
1
Expert's answer
2020-04-27T01:52:14-0400

"\Delta T=K_f m_l", ΔT - temperature change (T2-T1), Kf - the molal boiling point elevation constant, ml - molality or molal concentration of the solute in the solution.

n=m/M , n - chemical amount of substance(mol); m - mass of substance(g); M - molar mass of substance(g/mol).

M(C12H10O)=12g/mol*12+1g/mol*10+16g/mol=170g/mol

n(C12H10O)=17,98g/170g/mol=0,11mol

"m_l=n\/m" , n - chemical amount of substance (mol), m - solvent mass (CH3COOH)(kg)

291,7g=0,2917kg

ml=0,11/0,2917=0,38mol/kg

"\Delta T" =3,90oC/m*0,38mol/kg=1,48oC

T2=16,60oC

T1=T2-"\Delta T" =16,60-1,48oC=15,12oC

Answer:15,12oC.